OpenStudy (anonymous):

what are the intercepts of the equation 2x+3/2y+3z=6

OpenStudy (anonymous):

is it (3,0,0) (0,4,0)(0,0,2)

OpenStudy (amistre64):

2x+3 ------ -= 6 ; is this the equation? 2y+3z

OpenStudy (anonymous):

no its a fraction

OpenStudy (amistre64):

2x +3/2y +3z = 6 this then :)

OpenStudy (anonymous):

yes its a graphing equation

OpenStudy (amistre64):

we need to multiply everything by y; to get rid of that bit

OpenStudy (anonymous):

its gong to be 4

OpenStudy (amistre64):

2xy +3/2 +3zy = 6y is the sane equation, but has no limits when y = 0

OpenStudy (amistre64):

2xy +3/2 +3zy -6y = 0 y(2x +3z -6) +3/2 = 0 is another way of writing it

OpenStudy (amistre64):

when x and y = 0 there is no value..... is there?

OpenStudy (anonymous):

this is the answer i been got it (3,0,0) (0,4,0)(0,0,2)

OpenStudy (amistre64):

is that answer correct? or is that the answer you are checking? becasue y cant = 0 in this..

OpenStudy (amistre64):

(0,4,0) is good (x,0,0) is not valid, becasue y cant be 0 (0,0,z) is not valid cause y cant = 0

OpenStudy (amistre64):

in other words, yo have a function that only intersects the y axis; and never touches the other 2

OpenStudy (anonymous):

no, he means its a plane

OpenStudy (anonymous):

not a fraction

OpenStudy (anonymous):

people need to get the notation correct when asking a question , the fracts are coeffiecients of the plane

OpenStudy (amistre64):

the surface is a hyperbolic cylindar

OpenStudy (anonymous):

hello amistre!

OpenStudy (amistre64):

oh lol

OpenStudy (amistre64):

hi sat :)

OpenStudy (anonymous):

did you take yesterday off?

OpenStudy (amistre64):

i wish lol.... I spent the day under my front end fixing the steering parts...

OpenStudy (anonymous):

i meant from "off from here"

OpenStudy (anonymous):

and to the asker, yes the answers you posted in the first post are correct

OpenStudy (amistre64):

it was a forced hiatus lol

OpenStudy (anonymous):

oh yes. it is right. set two = 0 solve for third

OpenStudy (amistre64):

2x+(3/2)y+3z=6 4x +3y +6z = 12 x=3, y=4, z=2 (3,0,0) (0,4,0) (0,0,2) would then be the intercepts yes